\begin{thebibliography}{10}

\bibitem{AL}
Abresch, U.; Langer, J.:
{\sl The normalized curve shortening flow and homothetic solutions}. J. Differential Geom. 23(2): 175-196 (1986).

\bibitem{ang}
Angenent S.B.: {\sl Shrinking Doughnuts}. In: Lloyd N.G., Ni W.M., Peletier L.A., Serrin J. (eds) Nonlinear Diffusion Equations and Their Equilibrium States, 3. Progress in Nonlinear Differential Equations and Their Applications, vol 7. Birkhäuser, Boston, MA.

\bibitem{BBI}
Burago, D.; Burago, Y.; Ivanov, S.: 
{\sl A course in metric geometry}, Graduate Studies in Mathematics, Volume: 33; 2001

\bibitem{BW}
Bernstein, J. and Wang, L.: {\sl A Topological Property of Asymptotically Conical Self-Shrinkers of Small Entropy}. Duke Math. J. {\bf 166}, no. 3 (2017), 403-435.

\bibitem{yakov}
Berchenko-Kogan, Y: 
{\sl The entropy of the angenent torus is approximately 1.85122}. Experimental Mathematics, 0(0):1-8, 2019

\bibitem{BNS}
Buzano, R.; Nguyen, H.T.; Schulz, M.B.:
{\sl Noncompact self-shrinkers for mean curvature flow with arbitrary genus}. https://arxiv.org/abs/2110.06027

%\bibitem{CCMS}
%Chodosh, O.; Choi, K.; Mantoulidis, C.; Schulze, F.:
%{\sl Mean curvature flow with generic initial data}, https://arxiv.org/abs/2003.14344

\bibitem{CHH}
Choi, K.; Haslhofer, R. and Hershkovits, O.: {\sl Ancient low entropy flows, mean convex neighborhoods, and uniqueness} Acta Math. {\bf 228(1)} (2022), 217-301. 

\bibitem{CM1}
Colding, T.; Minicozzi, W.: {\sl Generic mean curvature flow I}. Ann. of Math. 175 (2012), 755-833

\bibitem{CM2}
Colding, T.; Minicozzi, W.: {\sl Smooth compactness of self-shrinkers}. Comment. Math. Helv. 87, No. 2, 463-475 (2012).

%\bibitem{DX}
%Ding, Q.; Xin, Y.L.:
%{\sl Volume growth eigenvalue and compactness for self-shrinkers}. Asian J. Math. 17, No. 3, 443-456 (2013).

\bibitem{CR}
Cecil, T. E., and Ryan, P. J.: Geometry of hypersurfaces (Vol. 10). New York: Springer (2015).
%\bibitem{DV}
%Dom{\' i}nquez-V{\' a}zquez, M.: {\sl An introduction to isoparametric foliations}, 2018.

%\bibitem{DLN}
%Drugan, G.; Lee, H.; Nguyen, X.H.
%{\sl A survey of closed self-shrinkers with symmetry}. Results Math. 73 (2018), no. 1, Paper No. 32, 32 pp.

\bibitem{DK}
Drugan, G., Kleene, S.J.: 
{\sl Immersed self-shrinkers}. Trans. Am. Math. Soc. 369, 7213-7250 (2017)

\bibitem{DN}
Drugan, G.; Nguyen, X. H.:
{\sl Shrinking doughnuts via variational methods}. J. Geom. Anal. 28 (2018), no. 4, 3725-3746.

%\bibitem{FCS}
%Fischer-Colbrie, D.; Schoen, R.:
%{\sl The structure of complete stable minimal surfaces of non-negative scalar curvature}. Commun. Pure Appl. Math. 33, 199-211 (1980).

\bibitem{HN}
Garcke, H.; Nürnberg, R.:
{\sl Numerical approximation of boundary value problems for curvature flow and elastic flow in Riemannian manifolds}. Numer. Math. 149, No. 2, 375-415 (2021).

%\bibitem{HW}
%Hershkovits, O.; White, B.:
%{\sl Sharp entropy bounds for self-shrinkers in mean curvature flow}. Geom. Topol. 23 (2019) 1611--1619.

\bibitem{Huisken}
Huisken, G.:
{\sl Asymptotic behavior for singularities of the mean curvature flow}. J. Differ. Geom. 31, No. 1, 285-299 (1990).

\bibitem{Ilmanen}
Ilmanen, T.:
{\sl Singularities of mean curvature flow of surfaces}. Preprint available at https://people.math.ethz.ch/~ilmanen/papers/sing.ps, 1995.

\bibitem{KKM}
Kapouleas, N.; Kleene, S. J.; M\o ller, N. M.:
{\sl Mean curvature self-shrinkers of high genus: non-compact examples}. J. Reine Angew. Math. 739, 1-39 (2018).

\bibitem{KMc}
Kapouleas, N. and McGrath, P.: {\sl Generalizing the Linearized Doubling approach, I: General theory and new
minimal surfaces and self-shrinkers}, preprint 2020, https://arxiv.org/abs/2001.04240.

%\bibitem{Ketover}
%Ketover, D.:
%{\sl Self-shrinking Platonic solids}. https://arxiv.org/abs/1602.07271

\bibitem{KM}
Kleene, S., M\o ller, N.M.:
{\sl Self-shrinkers with a rotational symmetry}, Trans. Amer. Math. Soc. 366 (2014), 3943-3963

\bibitem{MMM}
Ma, J.M.S.; Muhammad, A. M{\o}ller, N.M.: {\sl Entropy bounds, compactness and finiteness theorems for embedded self-shrinkers with rotational symmetry}. J. Reine Angew. Math. 793, 239-259 (2022).

\bibitem{MM}
Magni, A. and Mantegazza, C.: {\sl Some remarks on Huisken’s monotonicity formula for mean curvature flow}, in: Singularities in nonlinear evolution phenomena and applications, CRM Ser. 9, Edizioni della Normale, Pisa (2009), 157–169

\bibitem{Alex}
Mramor, A.: {\sl Compactness and finiteness theorems for rotationally symmetric self shrinkers}.
J. Geom. Anal. 31, 5094-5107 (2021)

\bibitem{AlexGT}
Mramor, A.:  {\sl On self shrinkers of medium entropy in $\mathbb R^4$}. https://arxiv.org/abs/2106.10243, to appear in Geom. Topol.

\bibitem{munzner}
M{\" u}nzner, H.F.: {\sl Isoparametrische Hyperfl{\" a}chen in Sph{\" a}ren}, Math. Ann. {\bf 251} (1980) no. 1 57-71

\bibitem{munzner2}
M{\" u}nzner, H.F.: {\sl Isoparametrische Hyperfl{\" a}chen in Sph{\" a}ren. II. {\" U}ber die Zerlegung der Sph{\" a}re in Ballb{\" u}ndel}, Math. Ann. {\bf 256} (1981), no. 2, 215–232.


\bibitem{Nguyen}
Nguyen, X. H.:
{\sl Construction of complete embedded self-similar surfaces under mean curvature flow, Part III}. Duke Math. J. 163 (2014), no. 11, 2023-2056.

\bibitem{Nicolaescu}
Nicolaescu, Liviu I. "The coarea formula." seminar notes. Citeseer. 2011.

\bibitem{PT}
Palais, R.S.; Terng, C.-L.:{\sl Reduction of variables for minimal submanifolds}, Proc. Amer. Math. Soc. {\bf 98} (1986), no. 3, 480–484.

\bibitem{riedler}
Riedler, O.: {\sl Closed embedded self-shrinkers of mean curvature flow}, J. Geom. Anal. 33, No. 6, Paper No. 172, 27 p. (2023).


%\bibitem{Klingenberg}
%Klingenberg, Wilhelm P.A.. {\sl Riemannian Geometry}, Berlin, New York: De Gruyter, 2011. https://doi.org/10.1515/9783110905120

%\bibitem{moller}
%M\o ller, N.M.:
%{\sl Closed self-shrinking surfaces in $\mathbb R^3$ via the torus}. arXiv:1111.7318

%\bibitem{Leonsimon}
%Simon, L.: 
%{\sl Asymptotics for a class of non-linear evolution equations, with applications to geometric problems}, Ann. Math. 118 (1983) 525-571


\bibitem{Stone}
Stone, A.:
{\sl A density function and the structure of singularities of the mean curvature flow}. Calc. Var. Partial Differ. Equ. 2, No. 4, 443-480 (1994).

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amsfonts}
\usepackage{latexsym,amsmath,amsthm,amssymb,amsfonts,mathrsfs}
\usepackage{graphicx}
\usepackage[usenames, dvipsnames]{color}
\usepackage{tikz}
\usepackage{float}
\usepackage{ulem}

\numberwithin{equation}{section}

\newtheorem{thm}{Theorem}[section]

\newtheorem{lem}[thm]{Lemma}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{dfn}[thm]{Definition}
\theoremstyle{remark}
\newtheorem{rem}{Remark}[section]

\newcommand{\pp}[1]{\frac{\partial}{\partial {#1}}}

\title{Entropy bounds for self-shrinkers with symmetries}
\author{John Man Shun Ma, Ali Muhammad}
\date{\today}

\begin{document}

\thanks{The authors were supported by the Danish National Research Foundation  CPH-GEOTOP-DNRF151.}


\begin{abstract}
In this work we derive explicit entropy bounds for two classes of closed self-shrinkers: the class of embedded closed self-shrinkers recently constructed in \cite{riedler} using isoparametric foliations of spheres, and the class of compact non-spherical immersed rotationally symmetric self-shrinkers. These bounds generalize the entropy bounds found in \cite{MMM} on the space of complete embedded rotationally symmetric self-shrinkers.
\end{abstract}

\maketitle

\section{Introduction}
A smooth hypersurface $\Sigma$ in $\mathbb R^{n+1}$ is called a {\sl self-shrinker} if 
\begin{equation}
H_{\Sigma}(x) = \frac{1}{2} \langle x, \nu\rangle, \ \ \text{ for all } x\in \Sigma.
\end{equation}
Here $\nu$ is the unit normal vector field on $\Sigma$, $H_\Sigma = \operatorname{div} \nu$ is the mean curvature of $\Sigma$ and $x$ is the position vector. Self-shrinkers are central objects of study in the mean curvature flow (MCF) since they serve as singularity models \cite{Huisken}, \cite{Ilmanen}. Besides the generalized cylinders, different techniques have been used to construct explicit examples of self-shrinkers \cite{AL}, \cite{ang}, \cite{KKM}, \cite{BNS}, \cite{Nguyen}, \cite{DN}, \cite{KMc}, \cite{riedler}.

In the seminal paper \cite{CM1} (see also \cite{MM}), Colding and Minicozzi defined the {\sl entropy} $\lambda(\Sigma)$ of any smooth hypersurface $\Sigma$. The entropy is a scaling and translation invariant quantity. More importantly, if $\{ \Sigma_t\}$ is a solution to the MCF, then $t\mapsto \lambda (\Sigma_t)$ is non-increasing and is constant if and only if (up to rescaling and space-time translation) $\Sigma_t = \sqrt{-t} \Sigma$ for $t \in (-\infty, 0)$, where $\Sigma$ is a self-shrinker. 

The entropy of the generalized cylinders were computed in \cite{Stone}, and the entropy of the Angenent doughnuts were numerically computed in \cite{yakov}, \cite{HN}. In general, upper bounds for the entropy are of great interest, since they normally imply surprising topological and geometrical consequences \cite{AlexGT}, \cite{BW}, \cite{CHH}, \cite{MMM}. 

In \cite{MMM}, the authors and M{\o}ller proved an entropy upper bound for complete embedded rotationally symmetric self-shrinkers in $\mathbb R^{n+1}$. Using that, several compactness and finiteness results for the space of such self-shrinkers were obtained. The goal of this paper is to generalize the entropy upper bounds obtained in \cite[Theorem 1.1]{MMM} in two directions. 

Using the theory of isoparametric foliations of spheres, Riedler constructed in \cite{riedler} new examples of closed embedded self-shrinkers in $\mathbb R^{n+1}$ with topology $\mathbb S^1 \times M$, where $M \subset \mathbb S^{n}$ is diffeomorphic to an isoparametric hypersurface of $\mathbb{S}^n$ for which the multiplicities of the principal curvatures agree. This is done by applying a reduction theorem from \cite{PT} to reduce the self-shrinker equation to a geodesic equation in an open subset of $\mathbb{R}^2$. Indeed, the construction generalizes the classical work by Angenent \cite{ang} for rotationally symmetric self-shrinkers. Our first result is the following theorem which provides entropy bounds for this class of self-shrinkers.

\begin{thm} \label{thm entropy upper bounds for f-invariant self-shrinkers}
Let $\mathscr M = \{ M_\varphi\}_{\varphi \in (0,\pi/g)}$ be an isoparametric foliation of $\mathbb S^n$ of type $(g, m,m)$, and let $N$ be a $f$-invariant embedded closed self-shrinker diffeomorphic to $\mathbb S^1 \times M$, where $M$ is diffeomorphic to a regular fiber of the foliation $\mathscr M $. Then there is a positive number $E_{g,m}$ such that $\lambda (N)\le E_{g,m}$. 
\end{thm}

We remark that the constants $E_{g,m}$ are explicit and depend only on the pair $(g, m)$. We refer the reader to section 3 for the terminology on isoparametric foliations appearing in Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers}.

Our second result is Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers} below which gives entropy bounds for immersed rotationally symmetric self-shrinkers diffeomorphic to $\mathbb S^1 \times \mathbb S^{n-1}$. Drugan and Kleene constructed in \cite{DK} infinitely many complete immersed rotationally symmetric self-shrinkers of various topological types. In the following theorem, we obtain an entropy upper bound for compact immersed non-spherical rotationally symmetric self-shrinkers in terms of the number of self-intersections of the corresponding profile curve. For each $n\ge 2$, let $E_n$ denote the entropy upper bound for complete embedded rotationally symmetric self-shrinkers in $\mathbb R^{n+1}$ obtained in \cite[equation (3.1)]{MMM}. 

\begin{thm} \label{thm entropy bounds for immersed rotationally symmetric self-shrinkers}
Let $\Sigma = \Sigma_\sigma$ be a compact immersed non-spherical rotationally symmetric self-shrinker in $\mathbb R^{n+1}$, where the profile curve $\sigma$ has $k$ self-intersection points counted with multiplicity. Then 
\begin{equation} \label{eqn entropy bound for immersed rotationally symmetric self-shrinkers}
    \lambda(\Sigma) \le (k+1) E_{n}. 
\end{equation}
\end{thm}

We refer the reader to section 4 and 5 for the terminology on rotationally symmetric self-shrinkers appearing in Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers}. Both Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers} and Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers} are proven using comparison theorems (\cite[Theorem 3.1]{MMM} and Theorem \ref{thm length bounds for immersed loop with k self intersections} respectively), which provide estimates for the length of closed geodesics $\sigma$ in a simply connected Riemannian surface with strictly positive Gaussian curvature. Theorem \ref{thm length bounds for immersed loop with k self intersections} also implies an entropy bound for a class of closed immersed $f$-invariant self-shrinkers (see remark \ref{remark immersed f-ivariant shrinkers}).



The organization of this paper is as follows. In section \ref{section intro to MCF}, we recall the definitions of self-shrinkers and the entropy of a hypersurface. In section \ref{section isoparametric foliation}, we recall the construction in \cite{riedler} and prove Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers}. In section \ref{section length bound on immersed curves}, we prove a general comparison theorem (Theorem \ref{thm length bounds for immersed loop with k self intersections}) and in the last section, we prove Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers}. 

The first author would like to thank Hojoo Lee for their discussion on immersed rotationally symmetric self-shrinkers. The authors would also like to thank Alexander Mramor for fruitful discussions. 

\section{Self-shrinkers and the entropy} \label{section intro to MCF}
Let $\Sigma$ be a properly immersed smooth hypersurface in $\mathbb R^{n+1}$. For any $x_0 \in \mathbb R^{n+1}$ and $t_0 >0$, the functional $F_{x_0, t_0}$ is defined in \cite{CM1} as 
\begin{equation}
F_{x_0, t_0} (\Sigma) = \frac{1}{(4\pi t_0)^{n/2}} \int_\Sigma e^{-\frac{\|x-x_0\|^2}{4t_0}} d\mu, 
\end{equation}
Where $d\mu$ is the volume form of $\Sigma$. It is well known (e.g. \cite[section 1]{CM2}) that the following statements are equivalent: 
\begin{itemize}
\item $\Sigma$ is a self-shrinker;
\item $\Sigma$ is a minimal hypersurface in $\mathbb R^{n+1}$ with respect to the conformal (Gau{\ss}ian) metric $g_{\text{\ss}}$, where 
\begin{equation} \label{eqn dfn of g_ang}
(g_{\text{\ss}})_{ij} = e^{-\frac{\|x\|^2}{2n}}\delta_{ij}
\end{equation}
and $\delta$ denotes the Euclidean metric on $\mathbb R^{n+1}$;
\item $\Sigma$ is a critical point of the $F$-functional $F = F_{0,1}$. 
\end{itemize}

The {\sl entropy} of a hypersurface $\Sigma$ is defined by (\cite{CM1}, \cite{MM})
\begin{equation}
\lambda (\Sigma) = \sup_{(x_0,t_0) \in \mathbb{R}^{n+1} \times \mathbb{R}_{>0}} F_{x_0, t_0} (\Sigma).
\end{equation}
We also need the following fact about the entropy of self-shrinkers \cite[section 7.2]{CM1}. 

\begin{lem} \label{lem lambda equals F 01}
Let $\Sigma$ be a properly immersed self-shrinker in $\mathbb R^{n+1}$. Then $\lambda (\Sigma) = F_{0,1} (\Sigma)$. 
\end{lem}

\section{Closed embedded self-shrinkers constructed by isoparametric foliations} \label{section isoparametric foliation}
In this section we prove Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers}. First, we recall the construction of the closed embedded self-shrinkers in \cite{riedler} using the theory of isoparametric foliations of spheres, which generalizes the classical construction of rotationally symmetric self-shrinkers by Angenent \cite{ang}. The definition of isoparametric foliations and the properties of such foliations of spheres, in particular, the structural theorems proved by M{\" u}nzner in \cite{munzner}, \cite{munzner2}, are neatly summarized in \cite[section 2]{riedler} and we closely follow their notation. For a general introduction, see for example \cite[Chapter 3]{CR}. We start by briefly recalling the necessary background on the theory. We then proceed to describe how the problem of finding self-shrinkers in the setting of \cite{riedler} is reduced to finding geodesics in a certain open subset of $\mathbb{R}^2$ equipped with a family of (incomplete) metrics $h_{g,m}=h$ with Gaussian curvature bounded from below by strictly positive numbers.

\subsection{Background on Isoparametric Foliations of $\mathbb{S}^n$} A smooth function $F: \mathbb{S}^n \to \mathbb{R}$ is called {\sl isoparametric} if there exist smooth functions $\phi_1, \phi_2 : F(\mathbb{S}^n) \to  \mathbb{R}$ such that 
\begin{align*}
    \|\text{grad} F\|^2 = \phi_1 \circ F, \quad \Delta F = \phi_2 \circ F,
\end{align*}
where $\text{grad} F$ and $\Delta F$ denote the gradient and the Laplacian of $F$, respectively. Foliations of $\mathbb{S}^n$ that arise from a family of level sets of an isoparametric function are called {\sl isoparametric foliations} of $\mathbb{S}^n$. Such level sets are then the fibers of a given foliation. By the work of M{\" u}nzner \cite{munzner}, the isoparametric function of a given foliation is a restriction to $\mathbb{S}^n$ of a homogeneous polynomial $F: \mathbb{R}^{n+1} \to \mathbb{R}$ of degree $g$ which satisfies certain differential equations, where $g\in \{1, 2, 3, 4, 6\}$. Such a polynomial $F$ is called the {\sl Cartan-M{\" u}nzner polynomial}. The foliation has exactly two singular fibers $V_\pm = (F|_{\mathbb S^n})^{-1} (\pm 1)$, and each regular fiber is given by 
\begin{equation}\label{definition regular fiber}
    M_\varphi = \{ x\in \mathbb S^n | \operatorname{dist}_0(x, V_-) = \varphi\},
\end{equation}
where $\varphi \in (0,\pi/g)$ and $\operatorname{dist}_0$ is the distance function on the round $\mathbb{S}^n$. On each $M_\varphi$, there are $g$ (the degree of the Cartan-M{\" u}nzner polynomial $F$) distinct principal curvatures, and they each assume a constant value on a given regular fiber. Furthermore, there are at most two distinct multiplicities for the principal curvatures which we denote by $m_1$ and $m_2$, and we have the following relation $n-1 = \frac{g}{2}(m_1+m_2)$. Also, for each $\varphi \in (0, \pi/g)$ we have:
\begin{equation}\label{regular fiber volume}
    \operatorname{Vol}_0 (M_\varphi) = c \sin \left(\frac{g}{2} \varphi \right)^{m_1} \cos \left(\frac{g}{2} \varphi \right)^{m_2},
\end{equation}
where $c$ is some constant independent of $\varphi$ and $\operatorname{Vol}_0 (M_\varphi)$ is the volume of $M_\varphi$ with respect to the round metric on $\mathbb{S}^n$.

We shall call an isoparametric foliation $\mathscr M = \{ M_\varphi\}$ of type $(g,m_1,m_2)$ if the fibers have $g$ distinct principal curvatures with corresponding multiplicities $m_1$ and $m_2$. In the following we shall restrict to the case where the multiplicities agree, i.e. $m_1=m_2$ and hence consider foliations of type $(g,m,m)$. In this case we have $n=mg+1$ and we shall use $n$ and $m$ interchangeably through this relation.


In the next lemma we determine the number $c$ in \eqref{regular fiber volume} in terms of $g$ and $m$. We start by recalling the following useful result: let $\pi : (E, g_E)\to (B, g_B)$ be a surjective Riemannian submersion between two Riemannian manifolds with compact fibers. 
Let $v : B \to \mathbb R$ be $v(x) = \operatorname{Vol}_{g_E} (\pi^{-1}(x))$. By the co-area formula \cite[Theorem 2.1]{Nicolaescu}, for all submanifolds $M$ of $B$ we have

\begin{equation} \label{eqn vol of preimage of M}
\operatorname{Vol}_{g_E} (\widetilde M) = \int_M v(x) d\mu_M (x),
\end{equation}
where $\widetilde M = \pi^{-1}(M)$ and $d\mu_M$ is the volume form of $M$ in $(B, g_B)$. 

\begin{lem} \label{lem volume of M_varphi}
Given an isoparametric foliation $\{ M_\varphi\}$ of $\mathbb S^n$ of type $(g, m, m)$, we have 
\begin{equation} \label{eqn volume of regular fiber of an isoparametric foliation}
\operatorname{Vol}_0(M_\varphi) = \frac{g\omega_n }{s(m)} \sin^m (g\varphi)
\end{equation}
for every regular fiber $M_\varphi$, $\varphi \in (0, \pi/g)$. Here $n = mg+1$,  $\omega_n$ is the volume of $\mathbb S^n$ and 
\begin{equation} \label{eqn dfn of s(m)}
s(m) = \int_0^\pi \sin ^m t \; \mathrm{d}t .
\end{equation}
\end{lem}

\begin{proof}
Since $m_1 = m_2 = m$ we have from \eqref{regular fiber volume} 
\begin{equation} \label{eqn vol of M_varphi in terms of c}
\operatorname{Vol}_0 (M_\varphi) = \frac{c}{2^m} \sin^m (g\varphi), 
\end{equation}
hence it suffices to find $c$. By the relation \eqref{definition regular fiber} we know that the mapping
\begin{align*}
    \mathbb S^n\setminus (V_- \cup V_+)&\to \left(0, \frac{\pi}{g}\right), \\
    x\in M_\varphi & \mapsto \varphi,
\end{align*}
is a Riemannian submersion, with respect to the standard metric on $\mathbb S^n$ and the Euclidean metric on the interval $(0,\pi/g)$. Using (\ref{eqn vol of preimage of M}) and (\ref{eqn vol of M_varphi in terms of c}), 
\begin{align*}
\omega_n := \operatorname{Vol}_0 (\mathbb S^n) = \int_0^{\pi/g} \operatorname{Vol}_0 (M_\varphi) d\varphi = \frac{c}{g2^m} \int_0^\pi \sin^m t \; \mathrm{d}t. 
\end{align*}
Hence $c = g2^m \omega_n s(m)^{-1}$ and Lemma \ref{lem volume of M_varphi} is proved.
\end{proof}
For simplicity, from now on we write 
\begin{equation} \label{eqn dfn of c_g,m}
c_{g, m} = \frac{g \omega_n}{s(m)}. 
\end{equation}
\subsection{The Reduction Theory; Proof of Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers}}
Equip $\mathbb R^{n+1} \backslash \mathbb R_{\ge 0} \cdot (V_+ \cup V_-)$ with the metric $g_{\text{\ss}}$ defined in (\ref{eqn dfn of g_ang}), and let 
\begin{align}\label{eqn definition of g_subm}
g_{\text{S}} = e^{-\frac{r^2}{2n}} (dr^2 + r^2 d\varphi^2)
\end{align}
be a metric on $(0, \infty) \times (0, \pi/g)$.
Define the map
\begin{align}
f : \mathbb R^{n+1} \backslash \mathbb R_{\ge 0} \cdot (V_+ \cup V_-) &\to (0,\infty) \times \left( 0, \frac{\pi}{g}\right),\\
\nonumber x &\mapsto \left( \| x\|, \frac{\arccos (F(x/\|x\|))}{g}\right). 
\end{align}
Here $F$ is the corresponding Cartan-M{\"u}nzner polynomial and $V_\pm = (F|_{\mathbb S^n})^{-1} (\pm 1)$ are the singular fibers of the foliation. It can be shown that $f$ is a Riemannian submersion. A set $N \subset \mathbb{R}^{n+1}$ is called $f$-{\sl invariant} if there exists a set $C \subset (0,\infty) \times (0, \pi/g)$ such that $N = f^{-1}(C)$. 

Using \cite[Theorem 4]{PT}, it is proved in \cite[Proposition 2.4]{riedler} that an $f$-invariant hypersurface $N \subset \mathbb{R}^{n+1}$ is a closed self-shrinker if and only if $C := f(N)$ is a closed geodesic in $(0,\infty) \times (0,\pi/g)$ with respect to the metric $h$ defined by
\begin{equation} \label{eqn definition of h}
h(r,\phi) := \operatorname{Vol}_{g_{\text{\ss}}} (f^{-1} (r, \varphi))^2 g_{\text{S}}(r,\phi).
\end{equation}
Note that for any $(r, \varphi) \in (0,\infty) \times (0,\pi/g)$, we have by Lemma \ref{lem volume of M_varphi}
\begin{align*}
\operatorname{Vol}_{g_{\text{\ss}}} (f^{-1} (r, \varphi)) & = \operatorname{Vol}_{g_{\text{\ss}}} (r M_\varphi) \\
&= e^{-\frac{(n-1)r^2}{4n}} \operatorname{Vol}_0 (r M_\varphi) \\
&= r^{n-1}e^{-\frac{(n-1)r^2}{4n}} \operatorname{Vol}_0 (M_\varphi) \\
&= c_{g,m} r^{n-1}e^{-\frac{(n-1)r^2}{4n}} \sin^m (g\varphi).
\end{align*}
Using \eqref{eqn definition of g_subm} we hence obtain the following expression for the metric $h$\footnote{In equation (4) of \cite{riedler}, the metric has the term $e^{-r^2}$ instead of $e^{-r^2/2}$ since they use a different scaling convention for self-shrinkers.}
\begin{equation}\label{eqn definition of h expression}
    h(r,\phi) =  c^2_{g,m}r^{2mg} e^{-r^2/2} \sin \left( g \varphi\right)^{2m}  (dr^2 + r^2 d\varphi^2).
\end{equation}
By a direct calculation, the Gaussian curvature $K = K_{g,m}$ of $h$ is given by
\begin{equation*}
K = \frac{ e^{r^2/2}}{c^2_{g,m} r^{2n} \sin \left( g \varphi\right)^{2m}} \left( r^2 + \frac{(n-1)g}{ \sin ^2 (g\varphi)} \right).
\end{equation*}
It is clear that $K$ is strictly positive and that $K \to +\infty$ as $(r, \varphi)$ tend to the boundary of $(0,\infty) \times (0,\pi/g)$. By simple calculus, the Gaussian curvature $K$ is bounded from below by $\kappa = \kappa_{g,m}>0$, where 
\begin{equation} \label{eqn expression for kappa_gm}
\kappa _{g,m}= \frac{ e^{\frac{y_{g,m}}{2}}}{c_{g,m}^2 y_{g,m}^n}\left( y_{g,m} + (n-1)g \right)
\end{equation}
and 
\begin{equation}
y_{g,m} =\frac{(n-1)(2-g) + \sqrt{(g-2)^2(n-1)^2 + 8ng(n-1)}}{2}.
\end{equation}
Moreover, by \eqref{eqn definition of h}, (\ref{eqn vol of preimage of M}) and \eqref{eqn dfn of g_ang},
the length $L_h (C)$ of the curve $C$ with respect to the metric $h$ is given by
\begin{align} \label{eqn length of C equal volume}
L_h (C) &= \int_C \operatorname{Vol}_{g_{\text{\ss}}} (f^{-1} (r, \varphi)) d\mu_C(r,\phi) \nonumber\\
&= \operatorname{Vol}_{g_{\text{\ss}}} (f^{-1} (C)) \nonumber \\
&= \operatorname{Vol}_{g_{\text{\ss}}} (N) \nonumber\\
&= \int_N e^{-\frac{\|x\|^2}{4}} d\mu_N. 
\end{align} 
Where $d\mu_C$ is the volume form of $C$ in $((0, \infty) \times (0, \pi/g), g_{\text{S}})$. 

\begin{proof}[Proof of Theorem \ref{thm entropy upper bounds for f-invariant self-shrinkers}]
For each pair $(g, m)$ let 
\begin{equation} \label{eqn expression for E_gm}
    E_{g,m} = \frac{2\pi }{(4\pi)^{n/2} \sqrt{\kappa_{g,m}}}. 
\end{equation} 
Let $N$ be a $f$-invariant embedded self-shrinker in $\mathbb R^{n+1}$. Then $C:= f(N)$ is the image of an embedded geodesic in $(0, \infty) \times (0, \pi/g)$ with metric $h$ given in (\ref{eqn definition of h}). Note that $N$ is diffeomorphic to $\mathbb S^1 \times M$ if and only if $f(N)$ is an embedded closed geodesic. Hence we can apply \cite[Theorem 3.1]{MMM} to conclude that 
$$L_h (C) \le \frac{2\pi }{\sqrt{\kappa_{g,m}}}.$$
Together with (\ref{eqn length of C equal volume}) and Lemma \ref{lem lambda equals F 01}, one obtains the entropy bound $\lambda (N)\le E_{g,m}$. 
\end{proof}

\begin{rem}
When $g=1$, the isoparametric foliation of $\mathbb S^n$ is given (up to congruence) by a family of $(n-1)$-spheres 
\begin{equation}
M_\varphi = \{ \cos\varphi \} \times \sin\varphi \ \mathbb S^{n-1}, \ \ \varphi \in (0,\pi) 
\end{equation}
and $f$-invariant hypersurfaces are percisely the rotationally symmetric hypersurfaces. In \cite{MMM}, the authors and M{\o}ller derived entropy bounds for complete embedded rotationally symmetric self-shrinkers. We remark that the entropy bounds found in \cite{MMM} are a special case of the ones obtained here: when $g=1$, then $m=n-1$ and we obtain from (\ref{eqn expression for E_gm}) that 
\begin{equation*}
E_{1, n-1} = \frac{2\pi}{(4\pi)^{n/2} \sqrt{\kappa_{1,n-1}}}. 
\end{equation*}
From (\ref{eqn expression for kappa_gm}) and $\kappa_n$ found in \cite[section 3]{MMM}, we have 
\begin{equation*}
\sqrt{ \kappa_{1,n-1}} = \frac{1}{c_{1,n-1}} \sqrt{\kappa_n}. 
\end{equation*}
Together with the identity $\omega_n = s(n-1) \omega_{n-1}$, (\ref{eqn dfn of c_g,m}) and the expression of $E_n$ in \cite[equation (3.1)]{MMM}, we have $E_{1,n-1} = E_n$ for all $n\ge 2$. 
\end{rem}

\section{Length upper bound on immersed closed geodesics in Riemannian surfaces with positive curvature} \label{section length bound on immersed curves}
In this section we prove Theorem \ref{thm length bounds for immersed loop with k self intersections}, which gives an upper bound on the length of closed immersed geodesics in a simply connected Riemannian surface with Gaussian curvature bounded below by a strictly positive constant $\kappa$. We start with a short outline of the section.

An immersed closed geodesic $\sigma$ encloses a number of domains. In Proposition \ref{prop gamma bounds k+1 domains} we determine the number of these domains in terms of the number of self-intersection points of $\sigma$. The proof of Theorem \ref{thm length bounds for immersed loop with k self intersections} then reduces to proving an estimate of the length of the boundary of each domain. This is done by showing that the domains are complete length spaces of curvature $\geq \kappa$, for which one can prove an analogous result as in \cite[Theorem 3.1]{MMM}. The necessary technical details are provided in Lemma \ref{shorest geo in Omega} and Proposition \ref{prop Omega domega has curvature ge kappa}. \\



Let $\mathbb V $ be a Riemannian $2$-manifold homeomorphic to $\mathbb R^2$. 

\begin{dfn}
Let $\gamma: \mathbb S^1 \to \mathbb V $ be an immersed closed curve. Let $\operatorname{Image}(\gamma)$ denote the image of $\gamma$. A point $p$ in $\operatorname{Image}(\gamma)$ is called a self-intersection point if $\gamma^{-1}(p)$ has more than one element. A self-intersection point $p$ is transverse if for all $t, s\in \gamma^{-1}(p)$, the pair $\gamma'(s), \gamma'(t)$ are not parallel to each other. Let $S$ be the set of all self-intersection points of $\gamma$ and let $k\in \mathbb N$. We say that $\gamma$ has $k$ self-intersection points counted with multiplicity if 
\begin{equation} \label{eqn dfn of k self intersection counted with multiplicity}
\sum_{p\in S} (|\gamma^{-1}(p)|-1) = k.
\end{equation}
\end{dfn}

\begin{dfn}
Let $\gamma$ be an immersed closed curve in $\mathbb V$ with only transverse self-intersections. A pre-compact connected component of $\mathbb V \setminus \operatorname{Image}(\gamma)$ is called a domain enclosed by $\gamma$.
\end{dfn}

\begin{prop}  \label{prop interior angles < pi}
Let $\gamma : \mathbb S^1 \to \mathbb V$ be an immersed closed curve in $\mathbb V$ with only transverse self-intersections. Then the boundary $\partial U$ of each domain $U$ enclosed by $\gamma$ is piece-wise smooth and at each corner of $\partial U$, the interior angle is less than $\pi$. 
\end{prop}

\begin{proof}
It is clear that the boundary of $U$ is piece-wise smooth. At each corner, the interior angle $\theta$ is not $\pi$ since $\gamma$ has transverse self-intersections. If $\theta >\pi$, then there is a sub-arc $\beta$ of $\gamma$ which lies in the interior of $U$ and $\beta$ is not part of $\partial U$. Let $\beta$ be the boundary of a domain $\tilde U$ enclosed by $\gamma$. But since $\beta$ lies in the interior of $U$, we must have $U = \tilde U$, which is a contradiction that $\beta$ is not part of $\partial U$. Hence we have $\theta < \pi$. 
\end{proof}

\begin{prop} \label{prop gamma bounds k+1 domains}
Let $\gamma$ be an immersed closed curve in $\mathbb V$ with $k$ transverse self-intersections counted with multiplicity. Then $\gamma$ encloses exactly $k+1$ domains.
\end{prop}

\begin{proof}
Let $U_1,\cdots, U_f$ be the domains enclosed by $\gamma$. For every $j=1, \cdots, f$ the closure of the domain $U_j$ has a piece-wise smooth boundary consisting of subarcs of $\gamma$. Let $\alpha_1 , \cdots, \alpha_e$ be the collection of those sub-arcs. Let $S = \{p_1, \cdots, p_\ell\}$ be the set of self-intersection points of $\gamma$ and let $W = \overline {U_1 \cup \cdots \cup U_f}$. Since $W$ is simply connected, the Euler formula implies 
\begin{equation}
\chi(W) = 1 = \ell - e + f,
\end{equation}
where $\chi(W)$ is the Euler characteristic of $W$. Let $s_i = |\gamma^{-1}(p_i)|$. By definition there are $2s_i$ sub-arcs with vertices $p_i$ (in the case where there is a loop, i.e., a sub-arc which starts and ends at $p_i$, that sub-arc would be counted twice). Since each sub-arc has exactly two vertices (also counted with multiplicity), we have
$$e = \sum_{i=1}^\ell s_i$$
and hence
\begin{equation}
f = 1 + e - \ell = 1 + \sum_{i=1}^\ell (s_i-1) = 1+k
\end{equation}
by (\ref{eqn dfn of k self intersection counted with multiplicity}). This finishes the proof of the proposition. 
\end{proof}

Let $h$ be a Riemannian metric on $\mathbb V$ such that the Gaussian curvature $K$ satisfies $K \geq \kappa > 0$ for some positive constant $\kappa$, and let $\sigma$ be an immersed closed geodesic in $(\mathbb V, h)$. By the uniqueness theorem of solutions to ODE, any self-intersection point of $\sigma$ must be transverse. We shall now describe the type of domains $U \subset \mathbb V$ that can be realized as the domains enclosed by $\sigma$. 

Let $U$ be a precompact, connected open subset in $(\mathbb V, h)$ and let $\Omega:=\overline U \subset \mathbb V$. Assume further that the boundary $\partial \Omega$ is parameterized by a curve $\beta : \mathbb S^1 \to (\mathbb V, h)$ consisting of piece-wise smooth geodesic arcs, and which furthermore satisfies the following properties: identifying $\mathbb S^1$ with $[0,d]$, with $0$ and $d$ identified, then there is a partition
$$\mathscr P =\{0 = t_0 < t_1 <\cdots < t_a = d (=0)\}$$
so that 
\begin{equation} \label{conditions on beta}
    \begin{cases} 
    \beta|_{\mathbb S^1 \setminus \mathscr P} \text{ is injective,} \\
    \text{for each }j=0, \cdots, a-1, \ \beta|_{[t_j, t_{j+1}]} \text{ is a geodesic in }(\mathbb V, h), \text{ and}\\
    \text{for each }j=0, \cdots, a-1 \text{ the interior angle at }q_j := \beta(t_j) \text{ is less than }\pi.
    \end{cases}
\end{equation} 

We remark that if $\beta$ is injective, then each vertex $q_j$ is joined by only two geodesic arcs $\beta|_{[t_{j-1}, t_j]}$, $\beta|_{[t_j, t_{j+1}]}$. 

Assume first that $\partial \Omega$ is parametrized by an injective curve $\beta$ which satisfies (\ref{conditions on beta}). Using Lemma \ref{shorest geo in Omega} below, we will define as in \cite[Section 3]{MMM} a metric structure $d^\Omega$ on $\Omega$ using the Riemannian metric $h$ and show that $(\Omega, d^\Omega)$ is a complete length space with curvature $\ge \kappa$, where $\kappa$ is the strictly positive lower bound of the Gaussian curvature of $(\mathbb V, h)$. See \cite{BBI} or \cite[Section 3]{MMM} for the terminology and notation needed from metric geometry. 

For each $p, q\in \Omega$, let 
$$ d^\Omega(p, q) = \inf_{\gamma} L_h(\gamma),$$
where the infimum is taken among piece-wise $C^1$ curves $\gamma : [0,1]\to \Omega$ with $\gamma(0) = p$, $\gamma(1) = q$ and 
\begin{equation}
    L_h(\gamma) =\int_0^1 \sqrt{h(\gamma'(t), \gamma'(t))} dt
\end{equation} 
is the length of $\gamma$ in $(\mathbb V, h)$. The following lemma is proved in \cite[Lemma 3.3]{MMM} when $\partial \Omega$ has no corners. 

\begin{lem} \label{shorest geo in Omega}
Let $\Omega$ be as above, i.e. a compact connected set in $(\mathbb V, h)$ so that the boundary $\partial \Omega$ is parametrized by an injective curve $\beta$ which satisfies (\ref{conditions on beta}). Let $p, q\in \Omega$. Then there is a simple geodesic $\gamma$ in $\Omega$ joining $p$ and $q$ which is shortest among all piece-wise $C^1$ curves in $\Omega$ joining $p$ and $q$. Moreover, 
\begin{itemize}
\item [(i)] $\gamma (t) \notin \{ q_1, \cdots, q_a\}$ for all $t\in (0,1)$.
\item [(ii)] if one of $p, q$ is in the interior of $\Omega$, then $\gamma$ also lies in the interior of $\Omega$ (except possibly at the other end point), 
\item [(iii)] if both $p, q$ are in $\operatorname{Image}(\beta) = \partial \Omega$, then either $\gamma$ lies completely in the image of $\beta|_{[t_j, t_{j+1}]}$ for some $j=0, \cdots, a-1$, or the interior of $\gamma$ lies inside the interior of $\Omega$.
\end{itemize}
\end{lem}

\begin{proof}
The Lemma follows from the same arguments used in the proof of \cite[Lemma 3.3]{MMM}. We only sketch the minor modifications here. 

Let $p, q\in \Omega$, $p\neq q$. As in the proof of \cite[Lemma 3.3]{MMM}, using a minimizing sequence we construct a Lipschitz length minimizing curve $\gamma : [0,1] \to \Omega$ with $\gamma(0)=p$, $\gamma(1) = q$, and we may assume that $\gamma ^{-1} (q) = \{1\}$, $\gamma^{-1}(p) = \{0\}$. 

First we show (i) by contradiction: if not, let $s_0 \in (0,1)$ be the first time such that $\gamma(s_0) = q_j$ for some $j$. Let $B$ be the closed geodesic ball of radius $\epsilon$ in $(\mathbb V, h)$ centered at $q_j$. For $\epsilon$ small enough, $p, q$ are not in $B$. Hence there are $s_1, s_2 \in (0,1)$ so that $s_1 < s_0< s_2$ and $\gamma (s_i) \in \partial B$ for $i=1,2$. Note that since $\beta|_{[t_{j-1}, t_j]}$ and $\beta_{[t_j, t_{j+1}]}$ are both geodesics, under geodesic polar coordinates centered at $q_j$, the set $B\cap \Omega$ is of the form
\begin{equation} \label{eqn U_x in polar coordinates}
    \{ (r, \theta) : 0\le r \le \epsilon, \theta_1 \le \theta\le \theta_2\}
\end{equation}
with $\theta_2 - \theta_1 <\pi$ by the third assumption in (\ref{conditions on beta}). Assuming $\epsilon$ is small enough so that $g$ is nearly Euclidean in $U$. Together with $\theta_2-\theta_1<\pi$, the (Euclidean) straight line $\ell$ joining $\gamma(s_1)$, $\gamma(s_2)$ has length
\begin{align*}
    L_h(\ell) < d_h (\gamma(s_1), q_j) + d_h (q_j , \gamma (s_2)) \le L_h (\gamma|_{[s_1, s_2]}). 
\end{align*}
Thus $\gamma$ cannot be length minimizing in $\Omega$ and hence (i) is proved. Part (ii) and (iii) of the Lemma can be proven similarly as in the proof of \cite[Lemma 3.3]{MMM} and are skipped. 
\end{proof}

From Lemma \ref{shorest geo in Omega} it follows that $(\Omega, d^\Omega)$ is a complete length space.

\begin{prop} \label{prop Omega domega has curvature ge kappa}
Let $\Omega$ be as in Lemma \ref{shorest geo in Omega}. Then the complete length space $(\Omega, d^\Omega)$ has curvature $\ge \kappa$, where $\kappa$ is the lower bound of the Gaussian curvature of $(\mathcal S, h)$. 
\end{prop}

\begin{proof}
Let $x\in \Omega$ and let $V_x$ be a geodesically convex neighborhood in $(\mathbb V, h)$ centered at $x$. When $x$ is in the interior of $\Omega$, we choose $V_x\subset \Omega$. Let $U_x = \Omega \cap V_x$. For any $p, q\in U_x$, let $\gamma$ be the unique shortest geodesic in $V_x$ joining $p$ and $q$. Arguing as in the proof of \cite[Proposition 3.6]{MMM}, it suffices to show that $\gamma$ lies in $U_x$. The argument there works in our case when $x\notin \{ q_1, \cdots, q_a\}$. For the case $x\in \{q_1, \cdots, q_a\}$, one may assume that $V_x$ is of the form (\ref{eqn U_x in polar coordinates}) for some $r_0 >0$ and $\theta_2 - \theta_1 <\pi$. Arguing as in the proof of \cite[Proposition 3.6]{MMM}, one shows that $\gamma$ lies in $H_{\theta_0}$ for all $\theta_0 \in [\theta_2, \theta_1 + \pi] $, where in geodesic polar coordinates  
$$H_{\theta_0} =\{ (r, \theta) : 0\le r\le r_0, \theta_0-\pi \le \theta\le \theta_0 \}.$$
Since 
$$U_x = \bigcap _{\theta_0 \in (\theta_2, \theta_1 + \pi)} H_{\theta_0},$$
one concludes that $\gamma$ lies in $U_x$. 
\end{proof}

Using exactly the same proof, by approximating the boundary of $\Omega$ by geodesic polygons, one has the following generalization of \cite[Theorem 3.1]{MMM}. 

\begin{thm} \label{thm length bound on boundary of Omega}
Let $\Omega$ be as in Lemma \ref{shorest geo in Omega}. Then the boundary $\partial \Omega$ satisfies 
\begin{equation}
|\partial \Omega|:= L_h(\beta) \le \frac{2\pi }{\sqrt{\kappa}}.  
\end{equation}
\end{thm}

Now we are ready to prove the main theorem in this section. 

\begin{thm} \label{thm length bounds for immersed loop with k self intersections}
Let $\sigma$ be an immersed geodesic loop in $(\mathbb V, h)$ with $k$ self-intersection points counted with multiplicity. Then the length $L_h(\sigma)$ of $\sigma$ satisfies
\begin{equation}
L_h(\sigma) \le (k+1) \frac{2\pi }{\sqrt{\kappa}}. 
\end{equation}
\end{thm}

\begin{proof}
By Proposition \ref{prop gamma bounds k+1 domains}, the immersed geodesic loop $\sigma$ encloses $k+1$ domains, denote the closure of these domains by $\Omega_1, \cdots, \Omega_{k+1}$. Fix each $i=1, \cdots, k+1$ and write $\Omega = \Omega_i$ for simplicity. The boundary $\partial \Omega$ of $\Omega$ is parametrized by a closed curve $\beta : \mathbb S^1 \to (\mathbb V, h)$ consisting of piece-wise smooth geodesic arcs which satisfies (\ref{conditions on beta}), but which may not be injective. Let $q$ be any vertex of $\partial \Omega$ (i.e. $q=\beta(t_j)$ for some $j=0, \cdots, a$). Let $s= |\beta^{-1}(q)|$. For any small $\epsilon >0$, let $B= B_\epsilon(q)$ be the closed geodesic ball in $(\mathbb V, h)$ centered at $q$ with radius $\epsilon$. Then $\beta$ intersects $\partial B$ at $2s$ points, given by $\exp_q(\epsilon e^{i \theta_0}), \cdots, \exp_q(\epsilon e^{i \theta_{2s-1}})$ in geodesic polar coordinates $(r, \theta)$ centered at $q$. Moreover, we have 
\begin{equation*}
\Omega \cap B = \{ \exp_q(r e^{i\theta}):  0\le r\le \epsilon, \ \  \theta_{2m}\le \theta \le \theta_{2m+1}, m=0, \cdots, s-1\}. 
\end{equation*}
For each $m$, let $\beta_m$ be the shortest geodesic in $B$ joining $\exp_q (\epsilon e^{i\theta_{2m}})$ and $\exp_q (\epsilon e^{i\theta_{2m+1}})$. Arguing as in the proof of Proposition \ref{prop Omega domega has curvature ge kappa}, $\beta_m$ also lies inside $\Omega$. Let $\Delta_m$ be the geodesic triangle in $\Omega$ with vertices $q$, $\exp_q (\epsilon e^{i\theta_{2m}})$ and $\exp_q(\epsilon e^{i\theta_{2m+1}})$. Let $D_\epsilon$ be the collection of all such triangles constructed for all vertices of $\partial \Omega$. Let $\Omega_\epsilon = \Omega \setminus D_\epsilon$. Note that the interior of $\Omega$ is homeomorphic to the interior of $\Omega_\epsilon$. Hence $\Omega_\epsilon$ is connected. Moreover, the boundary $\partial \Omega_\epsilon$ is parametrized by an injective curve consisting of piece-wise smooth geodesic arcs which satisfies (\ref{conditions on beta}). Thus we can apply Theorem \ref{thm length bound on boundary of Omega} to conclude 
\begin{equation*}
|\partial \Omega_\epsilon| \le \frac{2\pi}{\sqrt{\kappa}}. 
\end{equation*}
Taking $\epsilon \to 0$, we obtain 
\begin{equation*}
|\partial \Omega| \le \frac{2\pi}{\sqrt{\kappa}}
\end{equation*}
and hence 
\begin{equation*}
L_h(\sigma) \le \sum_{i=1}^{k+1} |\partial \Omega_i| \le (k+1) \frac{2\pi}{\sqrt{\kappa}}. 
\end{equation*}
\end{proof}

\section{Entropy bounds for immersed non-spherical closed self-shrinkers}
We recall that for any immersed curve $\sigma : I \to \mathbb H$, where $\mathbb H$ is the upper half space, the immersed rotationally symmetric hypersurface $\Sigma_\sigma$ in $\mathbb R^{n+1}$ with profile curve $\sigma (s)= (x(s), r(s))$ is given by 
\begin{equation*}
\Sigma_\sigma := \{ (x(s), r(s) \omega) : s\in I, \ \omega\in \mathbb S^n\}. 
\end{equation*}
Now we prove Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers}.

\begin{proof}[Proof of Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers}] By \cite[Proposition 2.3]{MMM}, $\Sigma_\sigma$ is a self-shrinker if and only if $\sigma$ is a geodesic in $(\mathbb H, g_A)$, where $g_A$ is given by 
\begin{align}\label{dfn of ang metric}
g_A = r^{2(n-1)} e^{-\frac{x^2 + r^2}{2}} (dx^2 + dr^2).
\end{align} 
Since $\Sigma_\sigma$ is compact, either $I = \mathbb S^1$ and $\Sigma_\sigma$ is an immersion from $\mathbb S^1 \times \mathbb S^{n-1}$. Or $I = (a, b)$, $\sigma (a^+), \sigma(b^-)$ lie in $\partial \mathbb H$ and $\Sigma_\sigma$ is an immersion from $\mathbb S^n$. Since $\Sigma_\sigma$ is non-spherical, the latter case is excluded and thus $\sigma$ is a closed immersed geodesic. By Theorem \ref{thm length bounds for immersed loop with k self intersections}, the length of $\sigma$ satisfies 
\begin{equation*}
L(\sigma) \le (k+1) \frac{2\pi}{\sqrt {\kappa_n}},
\end{equation*}
where $\kappa_n$ is the lower bound of the Gaussian curvature of $g_A$ computed in \cite[section 3]{MMM}. Together with
\begin{align} \label{relation between entropy and length}
 \lambda (\Sigma_\sigma) = (4\pi)^{-n/2} \omega_{n-1} L_A(\sigma)
\end{align}
and the numbers $E_n$ from \cite[equation (3.1)]{MMM}, given by
\begin{align*}
    E_n = \frac{2\pi \omega_{n-1}}{(4\pi)^{n/2} \sqrt{\kappa_n}}
\end{align*}
we obtain (\ref{eqn entropy bound for immersed rotationally symmetric self-shrinkers}). 
\end{proof}

\begin{rem}\label{remark immersed f-ivariant shrinkers}
    As in the proof of Theorem \ref{thm entropy bounds for immersed rotationally symmetric self-shrinkers}, one can use Theorem \ref{thm length bounds for immersed loop with k self intersections} to prove an entropy upper bound for the class of $f$-invariant closed self-shrinkers $N$ such that $f(N)$ is a closed immersed geodesic with $k$ self-intersections in $(0, \infty) \times (0, \pi/g)$ with respect to the metric $h$ given in \eqref{eqn definition of h expression}. However, as of now no such examples are known to the authors' knowledge.
\end{rem}

\bibliographystyle{abbrv}

\vspace{0.75cm}

    \noindent John Man Shun Ma \\
    Department of Mathematics, Southern University of Science and Technology \\
    No 1088, xueyuan Rd., Xili, Nanshan District, Shenzhen, Guangdong, China 518055 \\
    email: hunm@sustech.edu.cn

\bigskip

    \noindent Ali Muhammad \\
    Department of Mathematical Sciences, University of Copenhagen \\
    Universitetsparken 5, DK-2100 Copenhagen \O, Denmark\\
    email: ajhm@math.ku.dk

\end{document}